OpenStudy (anonymous):

Simplify the following expression. 5^0

zepdrix (zepdrix):

The exponent tells you how many 5's to multiply,\[\large\rm 5^1=5\]\[\large\rm 5^2=5\cdot5\]\[\large\rm 5^3=5\cdot5\cdot5\] \[\large\rm 5^0=\frac{5}{5}\]The zero power tells you that you're actually taking the 5 out of it.

OpenStudy (anonymous):

its 1 right

zepdrix (zepdrix):

Yes :)